Find the value of c that will make;
9x2+30x+c a perfect square trinomial
thank -you please help

Answer by stanbon(75887)

Find the value of c that will make;
9x2+30x+c a perfect square trinomial
-------------
Form:
(a+b)^2 = a^2+2ab+b^2
Your Problem:
a^2 = 9x^2 implies that a = 3x
2ab=30x implies that 2(3x)b = 30x implies that b = 5
---------------
Therefore 9x^2+30x+25 is a perfect trinomial square.
Therefore c = 25
===================
Cheers,
Stan H.